# Frequency-Explicit Shape Holomorphy in Uncertainty Quantification for Acoustic Scattering

### v1

@@ 8.2 Combined Smolyak-PML/FEM error bounds
With the Smolyak-quadrature error bound Thm. 8.2 in hand, we proceed to estimate the error in the quadrature-FE approximation of the corresponding mean of the far-field.
The integral (8.1a), is approximated by a Smolyak type quadrature formula $Q_{\Lambda}[\cdot]$ as in (8.6). We write, for any farfield direction ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$, and with the approximate farfield $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}$ as in (6.11),
$\mathbb{E}[\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}})]-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\\ =\mathbb{E}[\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}})-\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h})]+\mathbb{E}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\\ =:I+II.$ (8.9)
@@
Term $I$ is the expected discretization error of the PML-Galerkin FE discretization. For fixed direction ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$, estimated by the uniform (w.r.t. all admissible shapes $\breve{y}\in{\mathcal{P}}$) error bound in the computed farfield pattern $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}$ which was bound in Corollary 6.4, (6.12). Due to $\mu({\mathcal{P}})=1$,
@@
With $p\geq 1$, i.e. ${\mathbf{z}}\in{\cal H}_{p}$ as defined in (7.9), $\bar{p}$-summability of the weights $(\beta_{j})_{j\geq 1}\in(0,\infty)^{\mathbb{N}}$ satisfying (8.11) to bound term $I$ will imply, thanks to Theorem 8.2, quadrature convergence rates immune to the Curse of Dimensionality for Smolyak (or “sparse grid”) quadrature, as described in Section 8.1.2 and as analyzed in [66]. It will also imply, via the parametric $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy shown in corresponding results for higher-order Quasi-Monte Carlo (“HoQMC” for short) deterministic integration, as shown in [26, Thm. 3.1 and Prop. 4.1]. We leverage these results, with $k$-explicit error bounds, by verifying $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy (cp. Def. 8.1) of the parametric integrands ${\mathbf{z}}\mapsto\breve{u}^{\rm scat}_{\infty,\rm PML}({\mathbf{z}};\cdot)_{h}$.
@@
Then, there holds that the parametric PML-Galerkin FE farfield approximations
@@
This follows from [66, Lem. 3.3] with the summability exponent $\bar{p}$ bounded as in (8.14). ∎
With Prop. 8.3, the Smolyak error bound (8.8) then implies that with $p\geq 1$ as in (8.11)
@@
with the first term of the bound in turn majorized by (6.12). Absorbing the ($k$-independent) constants there, we arrive at the $k$-explicit error bound for the FE-PML and Smolyak-Quadrature approximated expected farfield pattern
$\left\|\mathbb{E}(\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}}))-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\right\|_{L^{2}(\mathbb{S}^{d-1};\mathbb{C})}\\ \lesssim\|\widehat{u}^{\mathrm{inc}}\|_{H^{1}_{k}(B_{2})}\left[\big{(}hk+(hk)^{p}k\big{)}(hk)^{p}k+e^{-C_{\rm PML,2}k\big{(}R_{\rm tr}-(1+\epsilon)R_{1}\big{)}}+k|{\mathrm{pts}}(\Lambda_{\epsilon},{\boldsymbol{\chi}})|^{-1-\frac{2p}{d-1}+\delta}\right]$ (8.16)
Here, the constant hidden in $\lesssim$ is bounded independent of $k$, but depends on the regularity $p\geq 1$ of the geometry, and on the space dimension $d=2,3$.
@@
In Prop. 8.3, we observed that the parametric integrands ${\mathbf{y}}\mapsto\breve{u}^{\rm scat}_{\infty,\rm PML}(\breve{{\mathbf{y}}};\cdot)_{h}$ are $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphic, uniformly with respect to $h$, as vector-valued integrand functions taking values in the (separable) Hilbert-space $X=L^{2}(\mathbb{S}^{d-1};\mathbb{C})$. For such integrand functions, certain deterministic so-called Quasi-Monte Carlo integration methods have been shown to furnish likewise higher orders of convergence, without the CoD. For general introduction and comprehensive presentation of this class of high-dimensional quadrature methods, we refer to [25].
Specifically, based on Prop. 8.3, it was shown in [26, Thm. 3.1, Prop. 4.1], that deterministic quadrature rules with $N$ points can be constructed that approximate the integral (8.1a) with rate $N^{-1/\bar{p}}$. The construction can be effected in $O(N\log N)$ many operations, subject to prior truncation to a suitable finite number $s$ of integration variables.
Particular QMC quadratures covered by the weighted norm setting in [26] allow for computable a-posteriori error estimation of the QMC integration error [27].

### v2

@@ 8.2 Combined Smolyak-PML/FEM error bounds
With the Smolyak-quadrature error bound Theorem 8.2 in hand, we proceed to estimate the error in the quadrature-FE approximation of the corresponding mean of the far-field.
The integral (8.1), is approximated by a Smolyak type quadrature formula $Q_{\Lambda}[\cdot]$ as in (8.6). We write, for any far-field direction ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$, and with the approximate far-field $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}$ as in (6.11),
[…]
@@
Term $I$ is the expected discretization error of the PML-Galerkin FE discretization. For fixed direction ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$, estimated by the uniform (w.r.t. all admissible shapes $\breve{y}\in{\mathcal{P}}$) error bound in the computed far-field pattern $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}$ which was bound in Corollary 6.4, (6.12). Due to $\mu({\mathcal{P}})=1$,
@@
With $p\geq 1$, i.e. ${\mathbf{z}}\in{\cal H}_{p}$ as defined in (7.9), $\bar{p}$-summability of the weights $(\beta_{j})_{j\geq 1}\in(0,\infty)^{\mathbb{N}}$ satisfying (8.11) to bound term $I$ will imply, thanks to Theorem 8.2, quadrature convergence rates immune to the Curse of Dimensionality for Smolyak (or “sparse grid”) quadrature, as described in Section 8.1 and as analyzed in [69]. It will also imply, via the parametric $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy shown in corresponding results for higher-order Quasi-Monte Carlo (“HoQMC” for short) deterministic integration, as shown in [28, Theorem 3.1 and Proposition 4.1]. We leverage these results, with $k$-explicit error bounds, by verifying $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy (cf. Definition 8.1) of the parametric integrands ${\mathbf{z}}\mapsto\breve{u}^{\rm scat}_{\infty,\rm PML}({\mathbf{z}};\cdot)_{h}$.
@@
Then, there holds that the parametric PML-Galerkin FE far-field approximations
@@
This follows from [69, Lemma 3.3] with the summability exponent $\bar{p}$ bounded as in (8.14). ∎
With Proposition 8.3, the Smolyak error bound (8.8) then implies that with $p\geq 1$ as in (8.11)
@@
with the first term of the bound in turn majorized by (6.12). Absorbing the ($k$-independent) constants yields an $k$-explicit error bound for the combined FE-PML, Smolyak-quadrature approximated expected far-field pattern.
Theorem 8.4.
Suppose that the dimension of the scatterer is $d\in\{2,3\}$ and that the assumptions of Theorem 5.8 hold. Assume further (8.11) for the weight sequence $\boldsymbol{\beta}=(\beta_{j})_{j\geq 1}$ ensuring that the scatterers’ geometry is $C^{p,1}$-regular, and the affine shape parametrization (7.4) with the expansion coefficients $\psi_{j}=\beta_{j}r_{j}$ with (7.2a), (7.2b) for $d=2,3$. Then the approximate, expected far-field pattern obtained by Smolyak-quadrature $Q_{\Lambda}$ in (8.6) applied to the parametric FE-PML approximations $\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}$ satisfies the error bound $\left\|\mathbb{E}(\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}}))-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\right\|_{L^{2}(\mathbb{S}^{d-1};\mathbb{C})}\\ \lesssim\|\widehat{u}^{\mathrm{inc}}\|_{H^{1}_{k}(B_{2})}\left[\big{(}hk+(hk)^{p}k\big{)}(hk)^{p}k+e^{-C_{\rm PML,2}k\big{(}R_{\rm tr}-(1+\epsilon)R_{1}\big{)}}+k|{\mathrm{pts}}(\Lambda_{\epsilon},{\boldsymbol{\chi}})|^{-1-\frac{2p}{d-1}+\delta}\right]\;.$ (8.16) Here, the constant hidden in $\lesssim$ is bounded independent of $k$, but depends on the regularity $p\geq 1$ of the geometry, and on the space dimension $d$.
@@
In Proposition 8.3, we observed that the parametric integrands ${\mathbf{y}}\mapsto\breve{u}^{\rm scat}_{\infty,\rm PML}(\breve{{\mathbf{y}}};\cdot)_{h}$ are $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphic, uniformly with respect to $h$, as vector-valued integrand functions taking values in the (separable) Hilbert-space $X=L^{2}(\mathbb{S}^{d-1};\mathbb{C})$. For such integrand functions, certain deterministic so-called Quasi-Monte Carlo integration methods have been shown to furnish likewise higher orders of convergence, without the CoD. For general introduction and comprehensive presentation of this class of high-dimensional quadrature methods, we refer to [27].
Specifically, based on Proposition 8.3, it was shown in [28, Theorem 3.1, Proposition 4.1], that deterministic quadrature rules with $N$ points can be constructed that approximate the integral (8.1) with rate $N^{-1/\bar{p}}$. The construction can be effected in $O(N\log N)$ many operations, subject to prior truncation to a suitable finite number $s$ of integration variables.
Particular QMC quadratures covered by the weighted norm setting in [28] allow for computable a-posteriori error estimation of the QMC integration error [29]. For related, recent work on wavenumber-explicit error estimates of QMC FE approximations in UQ, we refer to [38].
@@ 9 Conclusion
  • Finally, we note that we expect analogous shape holomorphy results to hold for the transmission problem for the time-harmonic Maxwell equations. Indeed, the Maxwell analogues of the Helmholtz bounds in [56] (used in §3) appear in [15], and the Maxwell analogue of the Helmholtz $h$-FEM convergence results of [35] (used in §5) appear in [12].
Appendix A Comparison of Corollary 6.4 with the results of [31]
Remark 6.5 described how the following two steps in the proof of Corollary 6.4 might appear over-simplistic:
  1. 1.
  2. 2.
Conversely, the paper [31] both
  1. 1.
    bounds the Galerkin error in terms of $\|f\|_{H^{s}}$ for arbitrary $s>0$, and
  2. 2.
A standard duality argument (see, e.g., [31, Proposition 2.1]) shows that
$\left\|{u-u_{h}}\right\|_{L^{2}}\leq C\left\|{u-u_{h}}\right\|_{H^{1}_{k}}\sup_{0\neq g\in L^{2}}\min_{v_{h}\in V_{h}}\frac{\left\|{\mathcal{S}^{*}g-v_{h}}\right\|_{H^{1}_{k}}}{\left\|{g}\right\|_{L^{2}}},$ (A.1)
and [31, Lemma 2.3] uses similar ideas to prove that
$\left|\int(u-u_{h})z\right|\leq C\left\|{u-u_{h}}\right\|_{H^{1}_{k}}\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}},$ (A.2)
If $z\in L^{2}$, then
$\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}}\leq C\Big{(}hk+(hk)^{p}k\Big{)}\left\|{z}\right\|_{L^{2}}$ (A.3)
by [54, 55], [14, Lemma 2.13], [35, Theorem 1.7]. Furthermore, if $z\in H^{p-1}$ is $k$-oscillatory, in that it satisfies the bound (5.14) (with $\widehat{f}$ replaced by $z$), then
$\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}}\leq C(hk)^{p}k\left\|{z}\right\|_{L^{2}}$ (A.4)
by [31, Part (ii) of Theorem 4.4]. The bounds (A.3) and (A.4) also hold with $\mathcal{S}^{*}$ replaced by $\mathcal{S}$ since $\mathcal{S}^{*}f=\overline{\mathcal{S}\overline{f}}$.
Therefore, if $u$ is the solution of the Helmholtz equation with $k$-oscillatory (in the sense of (5.14)) right-hand side $f\in H^{p-1}$ and $h$ is such that $(hk)^{p}k\leq C$ for sufficiently-small $C>0$, then the combination of (A.1), quasioptimality, (A.3), and (A.4) implies that
$\left\|{u-u_{h}}\right\|_{L^{2}}\leq C\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}k\left\|{f}\right\|_{L^{2}}.$ (A.5)
Furthermore, if $z\in H^{p-1}$ is $k$-oscillatory, as in the case of the far-field pattern (6.8), then the combination of (A.2), quasioptimality, and (A.4) implies that
$\left|\int(u-u_{h})z\right|\leq C(hk)^{p}k\left\|{f}\right\|_{L^{2}}(hk)^{p}k\left\|{z}\right\|_{L^{2}}.$ (A.6)
The bound (A.5) is the same as (6.14), and the bound (A.6) is no better than (A.5) when $\|z\|_{L^{2}}\sim 1$ and when $(hk)^{p}k$ is fixed (i.e., when one chooses the least restrictive condition on $h$ allowed by the theory as $k\to\infty$). That is, in this setting of data coming from a plane-wave with the quantity of interest being the far-field pattern expressed as the linear functional (6.8), the results of [31] indeed do not give a better result than Corollary 6.4. (Note that the bounds on the error in the functional in [31, Part (ii) of Corollary 4.5 and Part (ii) of Corollary 4.6] assume that $\|z\|_{H^{s^{\prime}}}$ is independent of $k$, and thus these results have better $k$-dependence than stated above.)